OpenStudy (anonymous):

If 5(2+4z)=85,what does z=?

OpenStudy (anonymous):

17=2+2z 15=4z z=15/4 z=3.75

OpenStudy (anonymous):

hint: 2 +4z = 17

OpenStudy (anonymous):

1)divide by 5 2)subtract 2 3)divide by 4 to get z